If 4+6+5=242038 and 3+8+9=242743 then what does 7+9+4= ?
Answer by Edwin McCravy(20054)

If 4+6+5=242038 and 3+8+9=242743 then what does 7+9+4= ?
This is an unusual use of an equal sign =, because those
are certainly not equal under the normal use of equal signs.
I will assume you mean that

4+6+5 corresponds to 242038 and
3+8+9 corresponds to 242743

And the question is,

What does 7+9+4 correspond to?

Now 4+6+5 = 15, which corresponds to 242038,
and 3+8+9 = 20, which corresponds to 242743.

So since 7+9+4 = 20 also, it must also correspond to the same number 
which 3+8+9 corresponds to.  So the answer is also 242743.

Edwin